# FeDaL: Federated Dataset Learning for Time Series Foundation Models

### v1

@@ 3.2 Global Bias Elimination
Motivation
While DBE effectively addresses spurious correlations at the client level in cross-domain time series, it remains insufficient for correcting global biases that emerge from heterogeneous temporal patterns and domain shifts across clients. For this, we propose Global Bias Elimination (GBE), a dual-component method that operates from the server side to mitigate representation divergence and improve global alignment. GBE comprises two components: (1) Representation Dynamic Correction, which corrects client-server gradient drift through a server-maintained correction term; and (2) Server-side Core-set Tuning, which refines the global model using privacy-preserving, client-selected data summaries. Together, these components address persistent cross-client discrepancies, allowing the global model to learn domain-invariant patterns while maintaining privacy.
Representation Dynamic Correction
Inspired by acar2021federated , we maintain a server-side state vector $\mathbf{s}$ that accumulates client-server gradient differences across communication rounds. This state serves as a correction term during global aggregation, helping to stabilize convergence and counteract representation drift. The state vector generation and global correction can be expressed as:
${\color[rgb]{0,0,1}\text{[State Updating] }}\mathbf{s}^{r}=\mathbf{s}^{r-1}-\beta\sum_{i}(\theta_{i}^{r}-\theta^{g,r-1})\quad{\color[rgb]{1,0,0}\text{[Global Correction]}}\quad\theta^{\prime\prime g,r}=\theta^{\prime g,r}-\frac{1}{\beta}\mathbf{s}^{r}$ (5)
Here, the initial state vector $s^{0}$ is initialized using the global model $\theta^{g,0}$, $\beta$ controls the strength of the correction, and $\theta^{\prime g,r}$ is the initial global model at the $r$-th round, obtained by weighted averaging mcmahan2017communication . By dynamically adjusting global updates in each round, this mechanism mitigates the impact of client-specific deviations and promotes a more generalized global representation.
Server-side Core-set Tuning
To further mitigate global bias, we introduce a server-side fine-tuning stage based on compact core-sets constructed locally by clients. Each core-set aims to approximate the client’s data distribution via gradient matching zhao2020dataset , selecting a small subset whose gradients resemble those of the full dataset. Given a local dataset $\mathcal{X}=\{x_{1},\dots,x_{Z}\}$, each client randomly samples a mini-batch $\tilde{\mathcal{X}}\subset\mathcal{X}$ of size $K\ll Z$ for efficiency and privacy. The objective is to construct an initial core-set $\mathcal{C}_{\text{init}}$ by minimizing the gradient discrepancy between $\mathcal{C}_{\text{init}}$ and $\tilde{\mathcal{X}}$:
$\mathcal{L}_{\text{match}}({\mathcal{C}}_{\text{init}})=\sum\nolimits_{x_{k}\in\tilde{\mathcal{X}}}\left|\nabla_{\theta}f_{\theta}(\mathcal{C}_{\text{init}})-\nabla\theta f_{\theta}(x_{k})\right|_{2}^{2},$ (6)
where $f_{\theta}$ denotes the current model on clients. The core-set is iteratively updated via: $\mathcal{C}_{\text{init}}\leftarrow\mathcal{C}_{\text{init}}-\eta\nabla_{\mathcal{C}_{\text{init}}}\mathcal{L}_{\text{match}}$, with learning rate $\eta$. To preserve privacy, we apply a Fourier-based perturbation mechanism that masks raw sequence details while retaining semantic structure. This is motivated by the observation that the phase of the Fourier transform encodes high-level temporal semantics, whereas the amplitude captures fine-grained variations. Accordingly, we perturb only the amplitude in the frequency domain: ${\mathcal{C}}^{\prime}={\mathcal{F}}^{-1}({\mathcal{F}}(\mathcal{C}_{\text{init}})+\epsilon{\mathcal{N}}(0,1))$, where $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the forward and inverse Fourier transforms, and $\epsilon$ controls the noise intensity. To ensure that $\mathcal{C}^{\prime}$ remains semantically faithful to the original data, we align its latent representation with that of the sampled mini-batch $\tilde{\mathcal{X}}$:
$\mathcal{L}_{\text{align}}({\mathcal{C}}^{\prime})=\frac{1}{K}\|\sum\nolimits_{c_{j}\in\mathcal{C}^{\prime}}f_{\theta}(c_{j})-\sum\nolimits_{x_{k}\in\tilde{\mathcal{X}}}f_{\theta}(x_{k})\|_{2}^{2}.$ (7)
The semantically aligned core-set $\mathcal{C}$, optimized from the perturbed $\mathcal{C}^{\prime}$ via Eq. 7, is then uploaded to the server. The server uses $\{\mathcal{C}_{i}\}^{N}$ to perform targeted fine-tuning on the global model, following the process of masked reconstruction nie2022time , yielding an refined version $\theta^{gt,r}$. To mitigate catastrophic forgetting and maintain continuity with prior global knowledge, we apply convex model fusion marfoq2022personalized :
$\theta^{g,r}=\alpha\theta^{\prime\prime g,r}+(1-\alpha)\theta^{gt,r},$ (8)
where $\alpha\in[0,1]$ is a fusion coefficient that controls the trade-off between stability and plasticity.
Algorithm 1 Implementation of our Federated Dataset Learning (FeDaL)
@@
1:Clients $\{c_{i}\}_{i=1}^{N}$, rounds $R$, local epochs $E$, core-set size $K$, period $\tau$; hyperparameters $\lambda$, $\alpha$, $\epsilon$.
2:Time Series Foundation Model $\theta^{g}=\{\theta^{b},\theta^{h}\}$
3:// Initialization Phase:
4:Server broadcasts $\theta^{g,0}=\{\theta^{b,0},\theta^{h,0}\}$ to initialize all clients; sets server state $\mathbf{s}=\mathbf{0}$.
5:Initialize global bias $\mathbf{b}^{g}$ by weighted averaging client biases $\hat{\mathbf{b}}^{p}_{i}$} after one warm-up epoch: $\mathbf{b}^{g}=:\sum_{i}\frac{n_{i}}{n}\hat{\mathbf{b}}^{p}_{i}$.
6:Client $i$ sets $\mathbf{b}_{i}^{\text{trend},0}=\mathbf{0}$, $\mathbf{b}_{i}^{\text{season},0}=\mathbf{0}$, and receives $\hat{\mathbf{b}}_{i}^{p,0}$.
7:for round $r=1$ to $R$ do
8:  Server samples client subset ${\mathcal{S}}_{r}$ with ratio $\rho$, and sends $\theta^{b,r-1},\theta^{h,r-1}$} to each client in ${\mathcal{S}}_{r}$
9:  // Local Updating Phase:
10:  for client $c_{i}\in\mathcal{S}_{r}$ in parallel do
11:    Obtain $\mathbf{h}_{i}^{\text{trend}},\mathbf{h}_{i}^{\text{season}}$ using time series decomposition $\triangleright$ Refer to Eq. 1
12:    Update local bias $\hat{\mathbf{b}}_{i}^{p,r}=\texttt{Mean}(\mathbf{h}_{i}^{\text{trend}})+\texttt{Mean}(\mathbf{h}_{i}^{\text{season}})$
13:    Optimize $\{\theta_{i}^{b,r},\theta_{i}^{h,r},\hat{\mathbf{b}}^{p,r}\}$ via SGD for $\min_{\theta_{i}}{\mathcal{L}}_{{\mathcal{D}}_{i}}(\theta_{i})$ with $\lambda$ for $E$ epoch $\triangleright$ Refer to Eq. 3
14:    Obtain core-set ${\mathcal{C}}_{i}$ via SGD for $\min{\mathcal{L}}_{\text{match}}({\mathcal{C}}_{\text{init}})$ and $\min{\mathcal{L}}_{\text{align}}({\mathcal{C}}^{\prime})$ with $K,\eta$ $\triangleright$ Refer to Eqs. 67
15:    Upload $\{\theta_{i}^{b,r},\theta_{i}^{h,r},{\mathcal{C}}_{i}\}$ to the server
16:  end for
17:  // Server Aggregation Phase:
18:  Update global model $\theta^{g,r}=\sum_{i}\frac{|\mathcal{D}_{i}|}{\sum_{j}|\mathcal{D}_{j}|}\theta_{i}^{r}$; Update server state $\mathbf{s}^{r}=\mathbf{s}^{r-1}-\alpha\sum_{i}(\theta_{i}^{r}-\theta^{g,r-1})$
19:  Obtain refined global model $\theta^{\text{gt},r}$ by fine-tuning on the collection of core-sets $(\bigcup_{i}\mathcal{C}_{i})$
20:  Update the TSFM $\theta^{g,r}=\alpha\cdot\theta^{\prime\prime g,r}+(1-\alpha)\cdot\theta^{\text{gt},r}$, Update global bias $\mathbf{b}^{g,r}=\frac{1}{|\mathcal{S}_{r}|}\sum_{i\in\mathcal{S}_{r}}\mathbf{b}^{p,r}_{i}$
21:end for
4 Experiments and Results
@@
4.1 Federated Time Series Representation Learning
@@
Learning dataset-invariant patterns from unlabeled and heterogeneous time series is critical for advancing TSFMs in federated settings. We evaluate FeDaL on two large-scale benchmarks: UTSD liu2024timer (7 domains, 1B) with domain-mixed (DM) partitioning, and CTSD chen2024federated (6 domains, 500M) with domain-independent (DI) partitioning. In DM, each client sees data from multiple domains, whereas in DI, clients are domain-specific. We compare FeDaL with five FL baselines, including FedAvg mcmahan2017communication , FedProx li2020federated , FedPer arivazhagan2019federated , FedRep collins2021exploiting , and TSFM-specific FFTS chen2024federated . All models are trained to reconstruct masked patches in an unsupervised manner nie2022time , using input length 1024, patch length 32, and batch size 1024. We use a masking ratio of 75% during training and report reconstruction MSE across varying mask ratios $\{20\%,35\%,50\%,75\%,90\%\}$ to assess robustness. The client join ratio $\rho=70\%$, with hyperparameters set as: core-set size $K=1024$, decomposition period $\tau=4$, and $\lambda=10^{-2}$, $\alpha=0.7$, and $\beta=10^{-1}$ for DBE and GBE. Each method is trained for 200 rounds. Full dataset statistics and implementation details are provided in Appendix˜B.
Table 1: Federated representation learning results under varying patch masking ratios. For UTSD, we simulate two levels of domain heterogeneity (H1 and H2) as detailed in the Appendix˜B. For CTSD, each dataset is treated as a domain-independent client. Bold indicates the best result, Underline the second best. $\dagger$} denotes evaluation using personalized models after server averaging; $\ddagger$} indicates client-side evaluation without aggregation.
Method
USTD Dataset (10 Clients, DM)
CTSD Dataset (18 Clients, DI)
20%
35%
50%
75%
90%
20%
35%
50%
75%
90%
H1 H2 H1 H2 H1 H2 H1 H2 H1 H2
@@
The main results of representation learning are presented in Table˜1, our FeDaL consistently outperforms baselines across varying heterogeneous settings. On USTD, it achieves superior results under both levels of domain heterogeneity (H1 and H2), and on CTSD, it excels across all masking ratios under domain-independent partitioning. Compared to the state-of-the-art federated TSFM pretraining method FFTS, FeDaL reduces reconstruction MSE by an average of 4.16% on USTD and 8.86% on CTSD, with particularly notable gains at higher masking ratios. These highlight our FeDaL’s ability to balance domain-specific feature learning on clients with domain-invariant aggregation at the server, leading to more generalizable representations under significant domain bias.
Figure 3: Results of ablation (Left) and hyperparameter sensitively (Right) on federated representation learning. Results on the UTSD dataset are averaged over H1 and H2. $\downarrow$ indicates performance drop; $\uparrow$, improvement relative to the original FeDaL. Experimental settings follow the main setup. Best viewed in color and with zoom.
Ablation and Hyperparameter Sensitively Results
We analyze the contribution of key components and the robustness of FeDaL to hyperparameter choices (Figure˜3). Ablations reveal that removing DBE or GBE significantly harms representation quality, confirming their role in addressing domain heterogeneity. Disabling DBE’s alignment step (Eq. 3) further degrades performance, emphasizing the importance of bias alignment. Removing Core-set or Correction (Eq. 5) impairs global adaptation, showing their necessity for model refinement under heterogeneity. Sensitivity results show: (i) overly large alignment weight $\lambda$ over-constrains local representations; (ii) larger core-sets (e.g., 2048, 3072) bring marginal gains at the cost of privacy and communication; (iii) extreme fusion weights $\alpha$ weaken generalization; (iv) unstable $\beta$ values impair client-server representation blending.
Bias Representation Analysis
Figure 4: Local bias change across communication rounds (R1 - R60) for select clients (Clients 3, 7, 9).
@@
To evaluate how effectively our FeDaL reduces domain bias during federated training, we visualize the evolution of local bias representations across three representative clients (Client 3, 7, 9) from round 1 to round 60, as shown in Figure˜4. At the initial round (R1), substantial variation is observed both across dimensions and among clients, reflecting strong intra- and inter-client biases due to limited generalization. As training progresses (e.g., R5 and R20), these biases progressively diminish. By round 60 (R60), the representations across clients exhibit strong alignment, indicating that FeDaL gradually encourages convergence toward domain-invariant features. This provides intuitive evidence that FeDaL effectively mitigates client-specific bias and facilitates the learning of generalized, cross-domain time series representations.
4.2 Downstream Task Adaption as Time Series Foundation Models

### v2

@@ 3.2 Global Bias Elimination
While DBE alleviates dataset-induced biases at clients, residual misalignment persists during aggregation, leading to divergence in global pattens. To mitigate this, we propose GBE with two components: (1) Gradient-level Dynamic Correction, which compensates for client-server drifts, and (2) Core-set Tuning, which refines the global model with privacy-preserving client summaries. GBE reduce cross-client discrepancies and guide the global model toward domain-invariant patterns.
Gradient-level Dynamic Correction.
The intuition is that naive aggregation (McMahan et al., 2017) assumes client updates are unbiased estimates of the global gradient. However, when client datasets are heterogeneous, this assumption fails, and the aggregated model drifts toward client-specific directions. Inspired by (Acar et al., 2021), we maintain a server-side state vector $\mathbf{s}_{r}$ that records the accumulated drift between client and global models across communication rounds:
$\mathbf{s}^{r}=\mathbf{s}^{r-1}-\beta\sum_{i}(\theta_{i}^{r}-\theta_{g}^{r-1}),$ (6)
where $\theta^{i}_{r}$ denotes the local model from client $i$ at round $r$, $\theta^{g}_{r-1}$ is the global model from the previous round, and $\beta$ is a scaling factor. This state serves as a correction memory: instead of directly adopting the naive FedAvg result $\tilde{\theta}_{g}^{r}$, we correct it as $\hat{\theta}_{g}^{r}=\tilde{\theta}_{g}^{r}-(1/\beta)\cdot\mathbf{s}^{r}$. Here, $\tilde{\theta}_{g}^{r}$ is the weighted average of client models, and $\hat{\theta}^{g}_{r}$ is the corrected global model used for the next round. By subtracting accumulated drifts, this mechanism prevents client-specific deviations (e.g., hospital-specific rhythms in healthcare or region-specific seasonality in energy data) from dominating the shared model.
Core-set Tuning
Even with gradient correction, global bias may persist if clients encode different structural patterns. To refine the global model further, we perform server-side fine-tuning using compact core-sets generated by clients. Each client first samples a small batch $\tilde{\mathcal{X}}\subset\mathcal{X}$ of size $K\ll|\mathcal{X}|$ for efficiency and privacy. From this, we construct an initial core-set $\mathcal{C}_{\text{init}}$, parameterized as a set of learnable vectors rather than raw samples, so that it can be directly optimized via gradient descent. To approximate the client-specific pattern, the core-set is optimized by minimizing gradient discrepancies against the sampled batch, using the gradient matching (Killamsetty et al., 2021) as:
[…]
Here the loss is well-defined because gradients are aggregated at the model-parameter level; the optimization updates $\mathcal{C}_{\text{init}}$ rather than $\theta$ (model parameters) with learning rate $\eta$. To protect raw data, we perturb $\mathcal{C}_{\text{init}}$ in the Fourier domain. Specifically, we perturb only the amplitude while keeping the phase, since the phase encodes semantic temporal patterns (e.g., periodicity), whereas the amplitude carries fine-grained details as $\mathcal{C}’=\mathcal{F}^{-1}(\mathcal{F}(\mathcal{C}_{\text{init}})+\epsilon\mathcal{N}(0,1))$, where $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the forward and inverse Fourier transforms, and $\epsilon$ controls the noise intensity. Perturbation may distort information, so we further align $\mathcal{C}$’ with the sampled batch $\tilde{\mathcal{X}}$ in latent space like Eq. 7:
[…]
The final aligned core-set $\mathcal{C}$ is uploaded to the server. The server fine-tunes the corrected global model $\hat{\theta}^{g,r}$ on aggregated $\mathcal{C}_{i}$} from clients, producing $\theta^{gt,r}$. To prevent catastrophic forgetting, we apply convex fusion as $\theta^{g,r}=\alpha\hat{\theta}^{g,r}+(1-\alpha)\theta^{gt,r}$, where $\alpha$ balances stability and adaptability.
FedDaL. As shown in Algorthm 1, FedDaL proceeds in three phases. In the initialization stage, the server broadcasts the model and establishes a global bias reference. In local updating, each client applies DBE to decompose features, estimate dataset-specific biases, and train the backbone with bias-regularized reconstruction, while also constructing a compact core-set for knowledge summarization. In the server aggregation stage, the server aggregates local models, applies dynamic correction strategy to counter client–server drifts, and refines the global model using compact core-sets. After each round, the server sends the updated model $\theta^{g}$ and the global bias $\mathbf{b}^{g}$ to selected clients, ensuring consistency across rounds. Through the interplay of DBE and GBE, our FedDaL progressively mitigates dataset-induced biases and aligns temporal patterns into a unified general TSFM.
Algorithm 1 Workflow of Federated Dataset Learning
@@
Clients $\{c_{i}\}_{i=1}^{N}$, total rounds $R$, core-set size $K$, decomposition period $\tau$, hyperparameters $\lambda$, $\alpha$, $\epsilon$
General TSFM $\theta^{g}=\{\theta_{b},\theta_{h}\}$
Initialization Phase:
Server initializes $\theta^{g,0}=\{\theta_{b}^{0},\theta_{h}^{0}\}$
Warm-up 1 epoch, broadcasts global bias $\mathbf{b}^{g,0}=\sum_{i}\tfrac{n_{i}}{n}\mathbf{b}_{i}^{0}$
Initialize server state $s=0$
for round $r=1$ to $R$ do
  Sample client subset $\mathcal{S}_{r}$ and send $\{\theta_{g}^{r-1},\mathbf{b}^{g,r-1}\}$
  Local Update (each $c_{i}\in\mathcal{S}_{r}$ in parallel):
  Decompose features into trend/seasonal terms (Eq. 1)
  Estimate local bias $\mathbf{b}_{i}^{r}=\operatorname{Mean}(\mathbf{h}_{t,i}^{r})+\operatorname{Mean}(\mathbf{h}_{s,i}^{r})$
  Train $\{\theta_{b,i}^{r},\theta_{h,i}^{r},\mathbf{b}_{i}^{r}\}$ by minimizing $\mathcal{L}_{i}$ (Eq. 4)
  Construct core-set $C_{i}^{r}$ (Eqs. 78)
  Upload $\{\theta_{b,i}^{r},\theta_{h,i}^{r},C_{i}^{r}\}$ to the server
  Server Aggregation:
  Aggregate models to obtain provisional global $\tilde{\theta}^{g,r}$
  Update server state $\mathbf{s}^{r}$ (Eq. 6)
  Apply gradient correction: $\hat{\theta}^{g,r}=\tilde{\theta}^{g,r}-(1/\beta)\cdot\mathbf{s}^{r}$
  Refine $\hat{\theta}^{g,r}\rightarrow\theta^{gt,r}$ by fine-tuning on core-sets
  Fuse models: $\theta^{g,r}=\alpha\hat{\theta}^{g,r}+(1-\alpha)\theta^{gt,r}$
  Update global bias $\mathbf{b}^{g,r}=\tfrac{1}{|\mathcal{S}_{r}|}\sum_{i\in\mathcal{S}_{r}}\mathbf{b}_{i}^{r}$
end for
4 Main Results
@@
4.1 Federated Representation Learning
@@
We evaluate FeDaL from an FL perspective, focusing on its ability to handle heterogeneous data. Experiments are conducted on two multi-domain datasets: UTSD (Liu et al., 2024c) with domain-mixed (DM) partitioning, where each client accesses data from multiple domains, and CTSD (Chen et al., 2025) with domain-independent (DI) partitioning, where each client is tied to a single domain. This allows us to examine FeDaL under both mixed-domain and domain-specific heterogeneity. We compare against five FL baselines: FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), FedPer (Arivazhagan et al., 2019), FedRep (Collins et al., 2021), and TSFM-oriented FFTS (Chen et al., 2025). All trained with patch-wise masked reconstruction (Nie et al., 2022) under 75% masking ratio with length of 512. Detailed implementation are provided in Appendix B.
@@
The main results are shown in Table 1, FeDaL consistently outperforms baselines across datasets and settings. Compared to the state-of-the-art federated TSFM pretraining method FFTS, FeDaL reduces reconstruction MSE by an average of 4.16% on USTD and 8.86% on CTSD. Notably, FeDaL adds about 2 MB per round due to transmitting core-sets, yet achieves significantly stronger generalization. These demonstrate FeDaL‘s ability to balance client-specific pattern learning with server-side alignment, yielding more transferable temporal patterns.
Bias Mitigation.
@@
To further assess its bias-mitigation capacity, we visualize the evolution of local bias features from clients in Figure 3. At R1, the biases show substantial variation across clients, reflecting strong intra- and inter-client discrepancies. As training progresses, these discrepancies shrink, then toward a common space. This provides intuitive evidence that FeDaL effectively reduces dataset-induced biases and promotes alignment toward domain-invariant features.
Table 1: Federated TS representation learning results averaged over masking ratios $\{20\%,35\%,50\%,75\%,90\%\}$. For UTSD, two heterogeneity levels (H1 and H2) are simulated as described in Appendix B, while in CTSD each dataset is treated as an independent client. Bold: the best, Underline: the second best. $\dagger$} indicates evaluation with personalized models after server averaging, and $\ddagger$} denotes client-side evaluation without aggregation. Full results in Table 16.
Method USTD (DM) CTSD (DI) Comm. Param\#
H1 H2
FedAvg 0.586 0.592 0.455 108.41 MB
FedProx 0.583 0.586 0.444 108.41 MB
FedPer 0.565 0.588 0.430 106.43 MB
FedRep 0.569 0.592 0.430 106.43 MB
FFTS 0.562 0.531 0.416 118.94 MB
Standalone 0.571 0.567 0.447
FeDaL (Ours) 0.551 0.511 0.387 110.41 MB
Figure 3: Local bias change across rounds (R1 - R60) for select clients (Clients 3, 7, 9).
Table 2: Ablation results. UTSD values are averaged over H1 and H2. “Vars.“ indicates the relative performance change compared with refees.
Model UTSD CTSD Average
MSE Vars. MSE Vars. MSE Vars.
FeDaL (Original) 0.573 0.405 0.489
w/o Alignment 0.602 $\downarrow$ 5.06% 0.434 $\downarrow$ 7.16% 0.518 $\downarrow$ 6.11%
w/o DBE 0.637 $\downarrow$ 11.17% 0.452 $\downarrow$ 7.16% 0.545 $\downarrow$ 9.17%
w/o Core-set Tuning 0.590 $\downarrow$ 2.97% 0.430 $\downarrow$ 6.17% 0.510 $\downarrow$ 4.57%
w/o Correction 0.600 $\downarrow$ 4.71% 0.431 $\downarrow$ 6.42% 0.516 $\downarrow$ 5.57%
w/o GBE 0.610 $\downarrow$ 6.46% 0.444 $\downarrow$ 9.63% 0.527 $\downarrow$ 8.05%
Ablation & Hyperparameter Sensitivity Results.
We analyze the contribution of five key components and the robustness of FeDaL to hyperparameter (Table 2). Ablations reveal that removing DBE or GBE significantly harms representation quality, confirming their role in addressing domain heterogeneity. Disabling DBE‘s alignment step (Eq. 4) further degrades performance, emphasizing the importance of bias alignment. Removing Core-set or Correction (Eq. 6) impairs global adaptation, showing their necessity for model refinement under heterogeneity. Sensitivity results (as shown in Figure 4) show: (i) overly large alignment weight $\lambda$ over-constrains local representations; (ii) larger core-sets bring marginal gains at the cost of privacy and communication; (iii) extreme fusion weights $\alpha$ weaken generalization; (iv) unstable $\beta$ values impair client-server representation blending.
Figure 4: Hyperparameter sensitivity. Results on UTSD are averaged over H1 and H2. $\downarrow$: performance drop; $\uparrow$: improvement relative to the original FeDaL. Best viewed in color and with zoom.
4.2 Adaption as General Time Series Foundation Models